5. Consider the following market demand for a good that consumers buy one unit at a time (e.g. smartphones, cars, game consoles, etc). Q=620-2P.

A market research shows that there are two groups of consumers, A and B, with different sensitivity to price change. Group A’s demand is estimated to be QA=200-0.5P and Group B’s demand is estimated to be QB=420-1.5P. Marginal Cost is assumed to be $30.

Given the information above, determine the most appropriate optimal pricing strategy and calculate the appropriate price(s). Show your answer(s) on a well-labeled graph(s).
Answer #1

Answert 5 Given thabi- | the following market demand fa a good that consumers bayone unft at a time. Erample: - (H) SmartphonQA = 200 - 0.51 00-p1-0-5 total deconne fd firm A=PQA = 2008-0-6p? MR(A)= dI RAldea a = 200-81-0.5 IMR(A)= 2p=400] a profit m200 280 310 Graph of green informaties